# LInK: Learning Joint Representations of Design and Performance Spaces through Contrastive Learning for Mechanism Synthesis

### v1

@@ Abstract
In this paper, we introduce LInK, a novel framework that integrates contrastive learning of performance and design space with optimization techniques for solving complex inverse problems in engineering design with discrete and continuous variables. We focus on the path synthesis problem for planar linkage mechanisms. By leveraging a multi-modal and transformation-invariant contrastive learning framework, LInK learns a joint representation that captures complex physics and design representations of mechanisms, enabling rapid retrieval from a vast dataset of over 10 million mechanisms. This approach improves precision through the warm start of a hierarchical unconstrained nonlinear optimization algorithm, combining the robustness of traditional optimization with the speed and adaptability of modern deep learning methods. Our results on an existing benchmark demonstrate that LInK outperforms existing methods with 28 times less error compared to a state-of-the-art approach while taking 20 times less time on an existing benchmark. Moreover, we introduce a significantly more challenging benchmark, named LINK-ABC, which involves synthesizing linkages that trace the trajectories of English capital alphabets—an inverse design benchmark task that existing methods struggle with due to large non-linearities and tiny feasible space. Our results demonstrate that LInK not only advances the field of mechanism design but also broadens the applicability of contrastive learning and optimization to other areas of engineering.
@@ 1 Introduction
Linkage mechanisms play a pivotal role in mechanical engineering, serving as fundamental components in applications ranging from the automation of manufacturing processes to the development of robotic systems. These mechanisms are crucial for translating input motions—often rotational—into desired output trajectories or paths, thereby enabling the execution of complex tasks through relatively simple inputs. The design of linkage mechanisms relies heavily on kinematics—the study of motion without considering the forces involved—and kinematic synthesis, the design process of mechanisms to achieve specific motions.
Despite the focused interest in kinematics and kinematic synthesis, designing complex kinematic systems remains a challenge Lipson (2008), often requiring trial and error, specialized expertise, or heuristic approaches to find effective solutions. The complexity arises because kinematic synthesis requires defining both the discrete components of a mechanism and their connections, as well as determining their continuous spatial locations, making it a mixed combinatorial and continuous problem.
In this paper, we focus on the inverse design of planar linkage mechanisms, where the design involves determining how many joints a mechanism has, what type of joints they are (either fixed or revolute joints, and how these joints are connected to one another using rigid linkages. On top of this, the initial position of the joints in space must also be determined. At each of these levels, mechanisms can be infeasible, and the majority of randomly sampled mechanisms tend to be invalid Heyrani Nobari et al. (2022). By examining the continuous values of initial positions, the non-linear and discontinuous nature of the problem becomes apparent. For instance, Figure 7 illustrates a mechanism that writes the letter “B”, where all design variables are held constant except for one joint (highlighted in red), which is moved in the x and y directions to demonstrate the feasible regions for these variables. The red-shaded area indicates where the mechanism fails to function. The three grey patches of feasibility for just one joint underscore the problem’s non-linearity and discontinuity, featuring multiple feasible regions. Furthermore, adjusting one joint alters the shape and number of feasible regions for all others. This discussion has focused solely on the continuous variables of joint positioning, yet the combinatorial challenge of defining the mechanism’s structure also plays a critical role. Not surprisingly, kinematic synthesis is sometimes regarded more as an art than a science due to its complexities 10. (2018).
@@
To demonstrate the effectiveness of our proposed framework, we focus on a specific type of kinematic synthesis, called, the path synthesis problem. This problem involves designing the aforementioned linkage mechanisms such that they can trace a desired curve with one circular motion of a single actuator. Planar linkage mechanisms use revolute and prismatic pairs to produce rotating, oscillating, or reciprocating motions which can combine into complex motions that trace desired curves. These mechanisms are commonly found in everyday machinery, such as in the moving parts of printing presses, auto engines, and robotics, making their design critical to a wide range of applications.
@@
To tackle these challenges of path synthesis in planar linkage mechanisms, we introduce Learning-accelerated Inverse Kinematics (LInK), which synergizes optimization with deep learning—termed “deep optimization”–to accelerate the process and surpass traditional methods in performance. Our approach utilizes a multi-modal, transformation-invariant contrastive learning framework to accurately represent and integrate the different modalities of performance and design spaces, enabling rapid retrieval of mechanisms from extensive datasets. Specifically, we train a multimodal contrastive learning model on a dataset of 10 million linkage mechanisms and their simulated kinematics, pre-compute embeddings for these mechanisms, and later retrieve them using embeddings of target curves produced by the same model (see Figure 3). Additionally, we refine these mechanisms through a hierarchical optimization algorithm, leveraging optimization’s inherent robustness and accuracy to improve the precision of our generated mechanisms. This integration of advanced engineering design techniques allows us to excel in designing linkage mechanisms, significantly surpassing existing methods in both efficiency and precision with 28 times better performance and a 20 times increase in speed compared to prior works on an established benchmark. We also create a significantly more challenging benchmark and show that our method can find solutions for all tasks in it too.
@@
  • We introduce a hierarchical deep optimization framework for integrating multi-modal contrastive learning-based warm-start with traditional optimization for inverse kinematic design of planar linkage mechanisms. We show that the proposed approach significantly outperforms the state-of-the-art in both speed and performance.
  • We develop a graph neural network framework named ’Hop Attention,’ designed to effectively capture features sensitive to the number of hops within a graph neural network. This feature is essential for analyzing planar linkage mechanisms.
@@ 2 Background and Related Works
2.1 Inverse Kinematics & The Path Synthesis Problem
Figure 2: Left: Demonstrates the path synthesis problem which involves a mixed combinatorial and conoth the combinatorial and continuous values. This figure also shows how mechanisms are represented as graphs. Right: This figure illustrates the path the solver takes to solve the kinematics of a given mechanism, showing how the skeleton of a mechanism (combinatorial design variables) is involved in the mechanism solution. Initially, the solver starts with the known joints (i.e., fixed and actuated joints highlighted red), and step by step the solver solves joints with two solved neighbors (Eqn. 10). In thtinuous nature and involves a non-linear solver whose solutions depend on bis example, three steps are taken to solve the mechanism. The numbered joints indicate the order of solution.
@@ 2.1 Inverse Kinematics & The Path Synthesis Problem
In our study of inverse kinematics, we focus on planar linkage mechanisms, which consist of interconnected rigid links joined by joints that move within a two-dimensional plane (see Fig. 2). These joints include revolute joints, which allow free rotation while keeping the linkage endpoints fixed, and fixed joints, which restrict the linkage end to a stationary position in 2D space. The mechanisms under study have a single degree of freedom, controlled by a rotary actuator connected to a fixed joint. This actuator’s movement dictates the entire mechanism’s configuration, determining the positions of all other joints and linkages based on its rotation.
@@
We model mechanisms as graphs. Let $M=(J,L)$ represent an undirected graph of a planar linkage mechanism, where $J=\{j_{1},j_{2},\ldots,j_{n}\}$ is the set of nodes (joints), and $L=\{(j_{i},j_{j})|j_{i},j_{j}\in J\}$ is the set of edges (linkages). Each node $j_{i}$ has a 2D position in space and a binary feature describing its type (fixed or free joint), denoted as $(x_{i},y_{i},T_{i})$, where $x_{i}$ and $y_{i}$ are the positions of joint $i$ in space and $T_{i}=1$ for fixed joints and $T_{i}=0$ for free joints (see Fig. 2). In our representation, joint $j_{0}$ is always fixed and joint $j_{1}$ is always free and there always exists a linkage $(0,1)\in L$ which connects joints $j_{0}$ and $j_{1}$ which is the driven linkage actuated around joint $j_{0}$ (see Fig. 2). In the specific problem we tackle in this paper, the goal is to generate a mechanism of any size ($n\geq 2$) such that the path traced by the final joint matches a given target path (see Fig. 2). This problem can be formulated as an optimization problem:
$\begin{split}\min_{M,\theta_{start},\theta_{end}}\quad&||F(M,\theta_{start},\theta_{end})-P_{target}||_{2}\\ \text{s.t.}\quad&DOF(M)=1\\ &\text{Singularity}(M,\theta_{start},\theta_{end})=0\end{split}$ (1)
Where $M$ is the graph describing the mechanism and $\theta_{start}$ and $\theta_{end}$ describe the range of motion for the actuator, $F(M)$ is the path traced by the mechanism $M$ and $P_{target}$ is the target path , while $DOF(M)$ is the degrees of freedom for mechanism $M$, which has to be always 1 (to be exactly defined by the motion of the single actuator), and $\text{Singularity}(M,\theta_{start},\theta_{end})$ is a function that determines if there are any singularities in a given mechanism $M$ for the actuator range of motion from $\theta_{start}$ to $\theta_{end}$.
@@
Singularities occur when the mechanism $M$ cannot be solved, leading to a locked state where the actuator cannot complete its intended range of motion (as shown in Figure 7 with the red region). The challenge involves defining the mechanism $M$, which includes specifying the number of joints (integer), their positions (continuous), their types (binary), and the linkage arrangement $L$. This arrangement $L$ can be represented by the adjacency matrix of $M$, containing $\frac{(n^{2}-n)}{2}$ binary values for $n$ joints. The problem is non-linear and involves different variable types and singularities are all non-linear and non-differentiable with respect to all or some of the optimization variables. Graph structure also adds combinatorial complexity. This complexity makes traditional optimization infeasible for large $n$, offering an opportunity for deep learning to address these difficult problems.
@@ 2.2 Computational Design of Planar Linkage Mechanisms
This section outlines significant research efforts aimed at solving inverse kinematics problems similar to the one discussed in this paper. The literature on path synthesis methods falls into three main categories: Numerical atlas-based, Optimization-based, and Deep learning-based approaches, with detailed discussions provided in Appendix A.
@@
Optimization-Based Approaches: These involve mostly modified versions of common conventional algorithms. Methods include genetic algorithms Lipson (2008); Khan et al. (2015b), Fourier descriptor-based optimization Ullah & Kota (1997); Wu et al. (2011), and mixed integer conic/non-linear programming (MICP, MINLP) Pan et al. (2023). These approaches often refine existing solutions Bächer et al. (2015); Thomaszewski et al. (2014) or are limited to specific problems Lipson (2008); Baskar & Plecnik (2021). Despite their generalizability, the effectiveness of these methods is limited by their computational intensity and they often struggle (as our results also show) to work for large mechanisms with many joints and target points due to the highly-constrained nature of the problem.
Deep Learning-Based Approaches: These aim to accelerate and enhance path synthesis by integrating traditional methods into data-driven frameworks. Techniques involve using generative models such as variational autoencoders (VAEs)Kingma & Welling (2014), clustering-based searchesDeshpande & Purwar (2019a; b; 2020), and reinforcement learning (RL) Fogelson et al. (2023). While promising, these methods either require retraining (RL-based methods) for new targets and can be limited by dataset size and mechanism type Deshpande & Purwar (2021); Vasiliu & Yannou (2001).
@@ 2.3 Contrastive Learning and Retrieval
Retrieval, in the context of machine learning, refers to the process of finding and returning relevant information from a dataset in response to a query. This can involve matching query features with features of items in the dataset to identify the best matches based on similarity metrics. While most deep learning retrieval research has focused on vision applications like content-based image retrieval (CIBR) Dubey (2022); Chen et al. (2023), our work shifts attention to cross-modal retrieval, particularly retrieving graphs of mechanisms based on their 2D kinematic paths. See Appendix B for more details.
Most cross-modal works involve retrieving images based on multi-modal information, such as text and other images. Early works like the correspondence autoencoder (Corr-AE) Feng et al. (2014) train autoencoders for text and images to create embeddings close to each other. The Deep Visual-Semantic Hashing (DVSH) model Cao et al. (2016) and similar methods like Textual-Visual Deep Binaries (TVDB) Shen et al. (2017) create joint embedding spaces for text-based retrieval. Adversarial learning approaches, such as Self-Supervised Adversarial Hashing (SSAH) Li et al. (2018), use adversarial networks to learn features and hash codes for different modalities Wang et al. (2017); Zhang et al. (2018); Gu et al. (2019); Ji et al. (2019). These methods primarily specialize in text and image, making them less adaptable to general cross-modal tasks. For our work, we need methods robust to general cross-modal retrieval, leading us to contrastive learning-based retrieval.
@@
$\mathcal{L}_{\text{CLIP}}$ $=-\log\frac{\exp(\text{sim}(f_{1},f_{2})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(f_{1},f_{j})/\tau)},$ (2)
where $\text{sim}(f_{i},f_{j})$ measures the cosine similarity between feature embeddings $f_{i}$ and $f_{j}$, and $\tau$ is the temperature parameter. This generalizability makes CLIP suitable for downstream tasks, such as multi-modal retrieval applications used in our methodology. Shared cross-modal embedding spaces have been shown to be effective for retrieval in such cases. For instance, Izacard et al. (2022) use contrastive embedding spaces for document retrieval. Similar approaches have been explored for tasks like temporal moment retrieval from videos based on text Zhang et al. (2021) and text-based molecular retrieval for molecule design Liu et al. (2023). Contrastive learning has immense potential for simultaneous modeling of design and performance spaces in engineering applications. For the current problem, we create a contrastive learning-enabled cross-modal approach for retrieving mechanisms based on their kinematic paths, with more details provided in subsequent sections.

### v2

@@ Abstract
In this paper, we introduce LInK, a novel framework that integrates contrastive learning of performance and design space with optimization techniques for solving complex inverse problems in engineering design with discrete and continuous variables. We focus on the path synthesis problem for planar linkage mechanisms. By leveraging a multimodal and transformation-invariant contrastive learning framework, LInK learns a joint representation that captures complex physics and design representations of mechanisms, enabling rapid retrieval from a vast dataset of over 10 million mechanisms. This approach improves precision through the warm start of a hierarchical unconstrained nonlinear optimization algorithm, combining the robustness of traditional optimization with the speed and adaptability of modern deep learning methods. Our results on an existing benchmark demonstrate that LInK outperforms existing methods with 28 times less error compared to a state-of-the-art approach while taking 20 times less time on an existing benchmark. Moreover, we introduce a significantly more challenging benchmark, named LINK-ABC, which involves synthesizing linkages that trace the trajectories of English capital alphabets—an inverse design benchmark task that existing methods struggle with due to large nonlinearities and tiny feasible space. Our results demonstrate that LInK not only advances the field of mechanism design but also broadens the applicability of contrastive learning and optimization to other areas of engineering. The code and data are publicly available at https://github.com/ahnobari/LInK.
@@ 1 Introduction
Linkage mechanisms play a pivotal role in mechanical engineering, serving as fundamental components in applications ranging from the automation of manufacturing processes to the development of robotic systems. These mechanisms are crucial for translating input motions—often rotational—into desired output trajectories or paths, thereby enabling the execution of complex tasks through relatively simple inputs. The design of linkage mechanisms relies heavily on kinematics the study of motion without considering the forces involved and kinematic synthesis, the design process of mechanisms to achieve specific motions.
Despite the reliance on the established practices of kinematics and kinematic synthesis, designing complex kinematic systems remains a challenge Lipson (2008). The design process often requires trial and error, specialized expertise, or heuristic approaches to find effective solutions. The complexity arises because kinematic synthesis requires defining both the discrete components of a mechanism and their connections, as well as determining their continuous spatial locations, making it a mixed combinatorial and continuous problem.
This paper focuses on the inverse design of planar linkage mechanisms, where the design involves determining how many joints a mechanism has, what type of joints they are (either fixed or revolute joints, and how these joints are connected to one another using rigid linkages. Further, the initial position of the joints in space must be determined. At each of these levels, mechanisms can be infeasible, and the majority of randomly sampled mechanisms tend to be invalid Heyrani Nobari et al. (2022). By examining the continuous values of initial positions, the nonlinear and discontinuous nature of the problem becomes apparent. For instance, Figure 3 illustrates a mechanism that writes the letter B, where all design variables are held constant except for one joint (highlighted in red), which is moved in the x and y directions to demonstrate the feasible regions for these variables. The red-shaded area indicates where the mechanism fails to function. The three grey patches of feasibility for just one joint demonstrate the problem’s nonlinearity and discontinuity, featuring multiple feasible regions. Furthermore, adjusting one joint alters the shape and number of feasible regions for all others. This discussion has focused solely on the continuous variables of joint positioning, yet the combinatorial challenge of defining the mechanism’s structure also plays a critical role. Not surprisingly, kinematic synthesis is sometimes regarded more as an art than a science due to its complexities 10. (2018).
@@
To demonstrate the effectiveness of our proposed framework, we focus on a specific type of kinematic synthesis, called, the path synthesis problem. This problem involves designing the aforementioned linkage mechanisms such that they can trace a desired curve with one circular motion of a single actuator. Planar linkage mechanisms use revolute and prismatic pairs to produce rotating, oscillating, or reciprocating motions which can combine into complex motions that trace desired curves. These mechanisms are commonly found in everyday machinery, such as in the moving parts of printing presses, auto engines, and robotics components, making their design critical to a wide range of applications.
@@
To tackle these challenges of path synthesis in planar linkage mechanisms, we introduce Learning-accelerated Inverse Kinematics (LInK), which synergizes optimization with deep learning—termed “deep optimization”–to accelerate the process and surpass traditional methods in performance. Our approach utilizes a multimodal, transformation-invariant contrastive learning framework to accurately represent and integrate the different modalities of performance and design spaces, enabling rapid retrieval of mechanisms from extensive datasets. Specifically, we train a multimodal contrastive learning model on a dataset of 10 million linkage mechanisms and their simulated kinematics, precompute embeddings for these mechanisms and later retrieve them using embeddings of target curves produced by the same model (see Figure 4). Additionally, we refine these mechanisms through a hierarchical optimization algorithm, leveraging optimization’s inherent robustness and accuracy to improve the precision of our generated mechanisms. This integration of advanced engineering design techniques allows us to excel in designing linkage mechanisms, significantly surpassing existing methods in both efficiency and precision with 28 times better performance and a 20 times increase in speed compared to prior works on an established benchmark. We also create a significantly more challenging benchmark and show that our method can find additional solutions within this new benchmark.
@@
  • We introduce a hierarchical deep optimization framework for integrating multimodal contrastive learning-based warm-start with traditional optimization for inverse kinematic design of planar linkage mechanisms. We show that the proposed approach significantly outperforms the state-of-the-art in both speed and performance.
  • We develop a graph neural network framework named ‘Hop Attention,’ designed to effectively capture features sensitive to the number of hops within a graph neural network. This feature is essential for analyzing planar linkage mechanisms.
@@ 2 Background and Related Works
2.1 Inverse Kinematics and The Path Synthesis Problem
Figure 2: This figure demonstrates the path synthesis problem which involves mixed combinatorial and continuous values. This figure also shows how mechanisms are represented as graphs.
@@ 2.1 Inverse Kinematics and The Path Synthesis Problem
In our study of inverse kinematics, we focus on planar linkage mechanisms, which consist of rigid interconnected links joined by joints that move within a two-dimensional plane (see Fig. 2). These joints include revolute joints, which allow free rotation while keeping the linkage endpoints fixed, and fixed joints, which restrict the linkage end to a stationary position in 2D space. The mechanisms under study have a single degree of freedom, controlled by a rotary actuator connected to a fixed joint. This actuator’s movement dictates the entire mechanism’s configuration, determining the positions of all other joints and linkages based on its rotation.
@@
We model mechanisms as graphs. Let $M=(J,L)$ represent an undirected graph of a planar linkage mechanism, where $J=\{j_{1},j_{2},\ldots,j_{n}\}$ is the set of nodes (joints), and $L=\{(j_{i},j_{j})|j_{i},j_{j}\in J\}$ is the set of edges (linkages). Each node $j_{i}$ has a 2D position in space and a binary feature describing its type (fixed or free joint), denoted as $(x_{i},y_{i},T_{i})$, where $x_{i}$ and $y_{i}$ are the positions of joint $i$ in space and $T_{i}=1$ for fixed joints and $T_{i}=0$ for free joints (see Fig. 2). In our representation, joint $j_{0}$ is always fixed and joint $j_{1}$ is always free. There always exists a linkage $(0,1)\in L$ which connects joints $j_{0}$ and $j_{1}$ which is the driven linkage actuated around joint $j_{0}$ (see Fig. 2). In the specific problem tackled in this paper, the goal is to generate a mechanism of any size ($n\geq 2$) such that the path traced by the final joint matches a given target path (see Fig. 2). This problem can be formulated as an optimization problem:
$\begin{split}\min_{M,\theta_{start},\theta_{end}}\quad&||F(M,\theta_{start},\theta_{end})-P_{target}||_{2}\\ \text{s.t.}\quad&DOF(M)=1\\ &\text{Singularity}(M,\theta_{start},\theta_{end})=0\quad.\end{split}$ (1)
Where $M$ is the graph describing the mechanism, and $\theta_{start}$ and $\theta_{end}$ describe the range of motion for the actuator. $F(M)$ is the path traced by the mechanism $M$, and $P_{target}$ is the target path. $DOF(M)$ is the degrees of freedom for mechanism $M$, which has to be always 1 (to be exactly defined by the motion of the single actuator), and $\text{Singularity}(M,\theta_{start},\theta_{end})$ is a function that determines if there are any singularities in a given mechanism $M$ for the actuator range of motion from $\theta_{start}$ to $\theta_{end}$.
@@
Figure 3: This figure shows the mechanism that traces B and the feasible region highlighted in blue, where the highlighted joint can move without causing a singularity. This shows the nonlinear and discontinuous nature of the problem and highlights the challenge of performing kinematic synthesis. This is only assuming all other joints keep the same positions. Changing the position of any joint could change the feasible region for all other joints making the problem very challenging from the perspective of conventional optimization.
Singularities occur when the mechanism $M$ cannot be solved, leading to a locked state where the actuator cannot complete its intended range of motion (as shown in Figure 3 with the red region). The challenge arises in defining the mechanism $M$, which includes specifying the number of joints (integer), their positions (continuous), their types (binary), and the linkage arrangement $L$. This arrangement $L$ can be represented by the adjacency matrix of $M$, containing $\frac{(n^{2}-n)}{2}$ binary values for $n$ joints. The problem is nonlinear and involves different variable types. Further, singularities are all nonlinear and non-differentiable with respect to all or some of the optimization variables, and graph structure adds combinatorial complexity. This complexity makes traditional optimization infeasible for large $n$, offering an opportunity for deep learning to address these difficult problems.
@@ 2.2 Computational Design of Planar Linkage Mechanisms
This section outlines significant research efforts aimed at solving inverse kinematics problems similar to the one discussed in this paper. The literature on path synthesis methods falls into three main categories: numerical atlas-based, optimization-based, and deep learning-based approaches. Detailed discussions are provided in Appendix A.
@@
Optimization-Based Approaches: These predominately involve modified versions of common conventional algorithms. Methods include genetic algorithms Lipson (2008); Khan et al. (2015b), Fourier descriptor-based optimization Ullah & Kota (1997); Wu et al. (2011), and Mixed-Integer Conic/Nonlinear Programming (MICP, MINLP) Pan et al. (2023). The approaches often refine existing solutions Bächer et al. (2015); Thomaszewski et al. (2014) or are limited to specific problems Lipson (2008); Baskar & Plecnik (2021). Despite their generalizability, the effectiveness of these methods is limited by their computational intensity and they often struggle (as our results show) to work for large mechanisms with many joints and target points due to the highly constrained nature of the problem.
Deep Learning-Based Approaches: These aim to accelerate and enhance path synthesis by integrating traditional methods into data-driven frameworks. The techniques involve using generative models such as Variational Autoencoders (VAEs)Kingma & Welling (2014), clustering-based searchesDeshpande & Purwar (2019a; b; 2020), and Reinforcement Learning (RL) Fogelson et al. (2023). While promising, these methods either require retraining (RL-based methods) for new targets or can be limited by dataset size and mechanism type Deshpande & Purwar (2021); Vasiliu & Yannou (2001).
@@ 2.3 Contrastive Learning and Retrieval
Retrieval, in the context of machine learning, refers to the process of finding and returning relevant information from a dataset in response to a query. This can involve matching query features with items’s features in the dataset to identify the best matches based on similarity metrics. While most deep learning retrieval research has focused on vision applications like Bontent-Based Image Retrieval (CIBR) Dubey (2022); Chen et al. (2023), our work shifts attention to cross-modal retrieval, particularly retrieving graphs of mechanisms based on their 2D kinematic paths. See Appendix B for more details.
Most cross-modal works involve retrieving images based on multimodal information, such as text and other images. Early works like the correspondence autoencoder (Corr-AE) Feng et al. (2014) train autoencoders for text and images to create embeddings close to each other. The Deep Visual-Semantic Hashing (DVSH) model Cao et al. (2016) and similar methods like Textual-Visual Deep Binaries (TVDB) Shen et al. (2017) create joint embedding spaces for text-based retrieval. Adversarial learning approaches, such as Self-Supervised Adversarial Hashing (SSAH) Li et al. (2018), use adversarial networks to learn features and hash codes for different modalities Wang et al. (2017); Zhang et al. (2018); Gu et al. (2019); Ji et al. (2019). These methods primarily specializing in text and image. This makes them less adaptable to general cross-modal tasks. Our work requires methods robust to general cross-modal retrieval, leading us to contrastive learning-based retrieval.
@@
$\mathcal{L}_{\text{CLIP}}$ $=-\frac{1}{N}\sum_{i=1}^{N}-\log\frac{\exp(\text{sim}(f_{i},g_{i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(f_{i},g_{j})/\tau)},$ (2)
where $\text{sim}(f_{i},g_{i})$ measures the cosine similarity between feature embeddings $f_{i}$ and $g_{i}$, and $\tau$ is the temperature parameter and $N$ is the number of samples (typically batch size during training). This generalizability makes CLIP suitable for downstream tasks, such as multimodal retrieval applications used in our methodology. Shared cross-modal embedding spaces have been shown to be effective for retrieval in such cases. For instance, Izacard et al. (2022) use contrastive embedding spaces for document retrieval. Similar approaches have been explored for tasks like temporal moment retrieval from videos based on text Zhang et al. (2021) and text-based molecular retrieval for molecule design Liu et al. (2023). Contrastive learning has immense potential for simultaneous modeling of design and performance spaces in engineering applications. For the current problem, we create a contrastive learning-enabled cross-modal approach for retrieving mechanisms based on their kinematic paths. More details are provided in subsequent sections.